(a) Calculate the nuclide masses of H (Z=1, A=1), H (Z=1,A=3), He (Z=2,A=4), Fe (Z=26, A=56),Ni(Z=28,A=57),Au(Z=79,A= 197),U(Z=92,A=235) using the Weizsacker Formula (where A=Z+N) for the binding energy of the nucleon - Use the coefficients (Mayo) provided in class – which are the 2 most dominant terms and which is the least dominant?

(b) Compare and plot/tabulate the predicted values with published values of the isotope masses and explain if the formula can be utilized for reasonably accurate (i.e., < 1% error) predictions all over the atomic mass space

(c) In which “A” region does the formula start to give unacceptable inaccuracies?
Answer #1

Weizsacker Formula for Binding energy B(A,Z) in MeV is given by

22) B(AZ) = QA - B.A2/3 - 72 41/3-SA, Z)

Where A is Mass number and Z is atomic number

A = Z+N , where N is number of neutrons

In the above formula a = 15.75 MeV , 8 = 17.8MeV , 7 = 23.7MeV ,

€ = 0.711 MeV,

11.18 73

\delta = 0 for odd-A, \delta=+1 for odd-odd nuclei and \delta=-1 for even-even nuclei

Following table lists the calculated Binding Energy value along with published results and % error

Element

A

Z

N

Calculated Binding Energy (MeV)

Calculated Binding Energy per nucleon(MeV)

Published value of Binding Energy per nucleon(MeV)

%error

H

1

1

0

-26.461

0

0

0

H

3

1

2

1.831528084

0.610509

2.827

78.5

He

4

2

2

22.35520289

5.588801

7.074

21

Fe

56

26

30

490.6614252

8.761811

8.79

0.3

Ni

57

28

29

488.8572616

8.576443

8.67

1.1

Au

197

79

118

1554.509533

7.890911

7.915

0.2

U

235

92

143

1785.898881

7.59957

7.59

0.1

For the region A \geq 56 , agreement of Binding energy per nucleon is \leq 1% error

Published values of Binding energy per nucleon for the above table is from https//periodictable.com